# Content-Style Disentangled Representation for Controllable Artistic Image Stylization and Generation

### v1

@@ 5.2.4 Quantitative Comparison
Tab. 1 presents the style similarity, image quality, text alignment and the overall subjective preference of our method compared with the state-of-the-art methods. We can see that our method achieves the highest style similarity and content alignment, demonstrating that our method, through decoupling representations, effectively fixed the problem of semantic conflict and captured the overall artistic style. Furthermore, users demonstrate a significantly greater preference for our method over other ones. More detailed results and explanations could be found in appendix Sec.A.3.
5.3 Content Style Decoupling
@@
To verify the decoupling effect, we conducted an experiment using portrait paintings as content and style respectively. As shown in Fig. 7, the generated images depict the content themes accurately: middle-aged men (first two rows), a young boy (third row), and middle-aged women (fourth and fifth rows). Vertically, the styles of the reference images are also well-preserved, with distinct brushstrokes in the third and fourth columns. This demonstrates the robustness of the proposed content-style disentanglement method.
5.4 Input Modality Ablation Study
To evaluate the effectiveness of multimodal representation training in the second phase, we conducted an ablation study. Table 2 shows the quantitative results under fixed combination and multimodal settings. Training with only specific content-style combinations leads to a noticeable drop in metrics, while exposure to diverse features improves generalization across styles and visual characteristics. This demonstrates that mixed training with multimodal content and style features is crucial for generating artistically expressive images that align with both content and style.
| Method | SS | IQ | TA |
| 0.286 | 5.694 | 0.304 | |
| multimodal | 0.293 | 5.811 | 0.308 |
6 Conclusion
In this paper, we proposed a method for disentangling content and style for artistic image stylization and generation. We constructed a multimodal art image-text dataset, WikiStyle+, to provide explicit data for supervised decoupling. We employed contrastive learning tasks to learn disentangled content and style representations, which then guided a diffusion model to generate stylized images. Our experiments across various tasks demonstrated the superiority of our method and highlighted the importance of effective content and style decoupling in image stylization.
Appendix A Appendix
A.1 Dataset Details
The word cloud map shows the distribution statistics of the WikiStyle+ dataset in three main categories: artists (such as Van Gogh, Monet), genres (posters, mosaics), and media (oil paintings, puzzles). In Figure 8, we present the selected statistics of dataset and its word cloud. The average length of style descriptions is 19 words, while content descriptions average 48 words. Style descriptions are more concise, focusing on key elements such as artist and the influence of specific style movements. In contrast, content descriptions tend to be more detailed, providing an overview of the artwork as a whole and often including more specific elements.
A.2 Model Details
We implemented our method based on Stable Diffusion v1.5. We employ the ViT-L/14 from CLIP (Radford et al., 2021) as the image encoder, maintaining 16 queries to extract style representations and 16 queries to extract content representations. Based on the pre-trained weight provided by the BERT model (Brooks et al., 2023), we initialize the Q-Former and then fine-tune the parameters of Query and Q-Former on WikiStyle+. In the second stage, in order to adapt to Stable Diffusion (Rombach et al., 2022), we adopted an IP-adapter architecture and froze the parameters of UNet (Ronneberger et al., 2015) and the text encoder, concentrating the training work only on the style module and two projection layers, rather than training the entire model from scratch. This targeted approach ensures that we can effectively fine-tune the model for specific tasks while leveraging the powerful foundation provided by Stable Diffusion. In the second stage of training, we adopt three modal tasks for joint training to enhance model generalization. We set the sampling rates of the three tasks to 1:1:1, and jointly train the style module, projection layer, and Q-Former.
A.3 Supplementary Experiments
A.3.1 Inference Speed
A.3.2 Quantitative comparison with SD and Artbank
In Table 4, we present quantitative comparisons with SD and ArtBank. It can be observed that our method achieves performance metrics comparable to Stable Diffusion and outperforms ArtBank. In our quantitative experiments for different settings, SD surpasses all existing methods. Combined with visual comparison in Figure 5, SD exhibits a more pronounced stylization, which contributes to its superior performance on quantitative metrics. However, as illustrated in Figure 16, SD demonstrates limitations in disentangled representations. The generated images sometimes include elements not present in the content image prompt but instead derived from common features of the original artwork.
A.3.3 Ablation study for components in disentangle loss(Eq(1))
In Eq.(1), Image-Text Contrastive Loss (ITC) primarily handles the tasks of alignment and disentanglement in the network, while the Image-Text Matching Loss (ITM) and Image-grounded Text Generation Loss (ITG) provide auxiliary support for modality alignment based on ITC. To evaluate the contribution of each loss function in training the multimodal alignment and disentanglement network, we conducted an ablation study. The results are shown in Figure 10 and Table 5, where four different configurations of the model are compared:
1) Training with only Image-Text Contrastive Loss (ITC). ITC is the core loss function in the first stage, enables the model to disentangle style and content more effectively by ensuring that features corresponding to style and content are aligned with their respective textual descriptions. When the CSDN is trained exclusively with ITC, the overall content structure is preserved. When the model is trained solely with ITC, the content alignment score achieves the highest value among all configurations. This indicates that as a contrastive loss, is highly effective in aligning the content features with their corresponding textual descriptions.
However, the generated results lack intricate brushstroke details, and fail to capture the characteristic yellowish tone of traditional oil paintings, as illustrated in Figure 10(a). The lack of fine-grained alignment with styles is corroborated by the decreased SS metrics in Table 5.
2) Training without Image-grounded Text Generation Loss (ITG): ITG trains the model to generate coherent style and content descriptions for a given image by predicting the next word or sentence based on the image and context. ITG ensures that the style and content information extracted from an image is not only disentangled but also interpretable and coherent, which enhances the overall textual-visual understanding of the model. As shown in Figure 10 (b), the absence of ITG affects the coherence of the style and content descriptions extracted from the image. For example, in the "A bus" prompt, the generated bus elements are inconsistent. Also, stylistic elements such as brushstrokes are less pronounced, leading to the drop of SS and TA in quantitative results.
3) Training without Image-Text Matching Loss (ITM): ITM operates as a binary classification task, predicting whether an image-text pair is a positive or negative match. This enables the model to focus on fine-grained correspondence between images and text, such as specific objects or elements mentioned in the prompts. When ITM is removed, the generated images show a loss of detail in both style and content alignment. For instance, in Figure 10 (c), the “house” element explicitly mentioned in the content prompt is missing in the generated image. This indicates that, without ITM, the model’s capacity to maintain fine-grained alignment is compromised, resulting in less coherent and contextually relevant outputs.
4) Full-loss setting: When all three loss functions are used together, the SS metrics achieve the highest scores. From Figure 10 (d), the generated images align closely with both style and content prompts, showcasing strong disentanglement and alignment. The style is faithfully preserved, while the content, such as the "house" or "bus," is accurately represented in the generated outputs. This demonstrates the complementary nature of the three loss functions in ensuring both disentanglement and multimodal coherence. While ITC prioritizes content alignment, the inclusion of ITM and ITG shifts the focus toward achieving a better balance, where both content and style are accurately disentangled and aligned with their textual descriptions.
| Method | SS | TA |
| 0.273 | 0.315 | |
| (w/o ) | 0.279 | 0.311 |
| (w/o ) | 0.278 | 0.313 |
| (full-loss) | 0.293 | 0.308 |
A.3.4 Style resemblance comparison with Dea-Diff and InstantStyle
In Figure 11, we provide a detailed comparison with InstantStyle and DEA-Diff, including the artist’s original painting in the last column to offer a more intuitive demonstration of how our results more closely follow the artist’s overall style. From the experimental results, InstantStyle focuses on replicating the colors and content of the reference image. DEA-Diff, on the other hand, faithfully reflects the content prompt but produces brushstrokes that deviate significantly from the reference image’s style, resembling comic illustrations rather than oil paintings. In contrast, our method emphasizes imitating the artist’s overall style, delivering more cohesive and stylistically accurate results. This is because DEA-Diff and InstantStyle define style differently from our approach. They focus on imitating the color and texture, whereas our method aims to replicate the artist’s overall painting habits.
Under a unified style, our approach reflects different content in varied yet stylistically consistent ways. From Figure 11, regardless of the input content, DEA-Diff and InstantStyle consistently produce the same color palette and textures, which do not align with the artist’s actual painting habits. On the contrary, our generated results incorporate similar colors and distinctive swirl patterns from Starry Night only when the input content prompt explicitly includes "starry night." Additionally, when the content prompt specifies a "city skyline," our results accurately depict urban scenes, while InstantStyle continues to replicate the rural content from the reference image. Meanwhile, DEA-Diff produces results that resemble anime rather than an oil painting. These observations demonstrate that our method effectively disentangles style and content, producing outputs that align more closely with the artist’s painting habit and overall style.
A.3.5 Impact of Content and Style Detail Levels on Disentanglement Performance
@@
From subfigure (a), we observe that as the textual descriptions of content progress from simple to complex, the generated images faithfully reflect the content information described in the text. When the content is described through an artistic image, our method perfectly disentangles the content information from images, accurately capturing elements such as the flowers, vase, table, and teacup. Meanwhile, the generated image consistently mimics Monet’s brushstrokes and artistic style, regardless of the level of detail in the content description.
@@
A.3.6 Comparison with StyleDrop
In Figure 13, we present comparisons with StyleDrop Sohn et al. (2023). For fair comparison, we directly used the original images from their paper. The experimental results show that StyleDrop and our method have different focuses. StyleDrop fine-tunes a single reference image and content prompt, with the generated results primarily transferring the colors and textures of the reference image. In contrast, our method leverages a disentanglement design to extract style information from the reference image, capturing the artist’s brushstroke techniques and overall style. This allows our approach to generate diverse outputs that align with the artist’s overall style when combined with different content prompts. Additionally, we include original works by the reference artist (Van Gogh) on similar themes to the content prompts, providing a clearer demonstration that our results better follow the artist’s brushstrokes and color patterns.
A.3.7 Comparison with DreamStyler
In Figure 14, we compare our method with DreamStyler Ahn et al. (2024). From the results, we observe major content leakage from the reference images in the first and second rows, where the generated results almost replicate the structure of the reference images but fail to adequately reflect the information provided by the content prompts. In the fourth row, the generated result resembles a photograph, failing to follow the style of the reference image. This limitation arises because DreamStyler focuses on enhancing content information in the generated results through text inversion during fine-tuning. As a result, when there is a significant mismatch between the content prompt and the reference style image, the disentanglement performance lacks robustness. In contrast, our method focuses on the explicit disentanglement of style and content information, achieved during the first stage of training. This approach ensures that the generated results not only faithfully adhere to the input content prompt but also closely align with the artist’s overall style.
A.3.8 Disentangle performance comparison with Dea-Diff
Additionally, Fig. 15 presents a visual comparison between our method and DeaDiff. We utilize images as content prompts and various style attributes as style prompts. Both our method and DeaDiff accurately capture the theme of the content image. However, DeaDiff exhibits several failure cases in disentangle style and content information from images. For example, under the "pencil" style, the color of the mountains in the content image is incorrectly treated as style information and transferred to the generated result. With "Samuel Peploe" style, DeaDiff struggles to produce a stylized output, instead following the photographic style from the content image. These issues arise because Dea-Diff lacks the explicit disentanglement design proposed in our method, resulting in suboptimal disentanglement capability. Consequently, it often reproduces the style of the content image or the content of the style image in the generated results. In contrast, our method achieved robust disentanglement across scenarios where content and style inputs are provided as either images or text, while faithfully follow the artist’s style.
A.3.9 Image style transfer
Additionally, although our model employs a non-reconstructive learning method and does not impose constraints for structural preservation, we adopt ControlNet (SoftEdge) to achieve image-based stylization with spatial control and compared our results with existing image style transfer methods: CycleGAN (Zhu et al., 2020), F-LSeSim (Zheng et al., 2021),
Fig. 16 illustrates visual comparison with structure-preserved image style transfer. We can see that with spatial control, our method can achieve structure-preserved image style transfer. Although our model employs a non-reconstructive learning method, the generated results successfully transfer styles from reference image while maintaining the structure of the content image, avoiding the introduction of elements from the original collection that do not exist in the content image, such as the results in the fourth row from SD, InST, and Artbank.
Table 6 presents the style similarity, image quality and the overall subjective preference of our method compared with the state-of-the-art methods. We can see that our method achieves the comparable style similarity and content alignment performance. This is understandable since our method is not specifically designed for structure preservation, which would inevitably affect the visual effect of the generated results. Still, users demonstrate a significantly greater preference for our method over other ones. More detailed results and explanations could be found in appendix Sec.A.3.
| Metrics | CycleGan | CUT | F-LSeSim | DiffuseIT | SD | InsT | ArtBank | Ours |
| SS | 0.263 | 0.266 | 0.237 | 0.218 | 0.293 | 0.296 | 0.288 | 0.283 |
| IQ | 5.407 | 4.902 | 5.263 | 5.081 | 5.626 | 5.630 | 5.470 | 5.506 |
| SP | 2.778 | 3.000 | 2.556 | 2.556 | 2.833 | 2.333 | 2.761 | 3.857 |
A.3.10 User Study
Aesthetic perception to human observers is often the ultimate goal for style transfer tasks. We conduct user study to evaluate the proposed algorithm against the state-of-the-art methods. We use the most commonly mentioned 7 styles: Vincent van Gogh, Claude Monet , Paul Cezanne , Samuel Peploe , Berthe Morisot ,Paul Gauguin and Ukiyo-e. For each comparison method, we synthesized 30 images, and obtained 390 images in total. For each participants, we randomly select 35 content-style-result triplets, and display them in random order. We ask participants to rate the images to gauge their preference for these images. The scoring scale ranges from 1 to 5, where 1 represents the least liked and 5 represents the most liked. A total of 40 participants took part in our survey, rating 35 images based on their preferences.
A.3.11 Content Style Disentanglement
To verify the decoupling effect, we conducted an experiment where we use still life paintings as content and style respectively. As shown in Fig. 17, when we observe horizontally, we can see that the generated images successfully depict the themes of the content images: the first two rows features yellow flowers, the third-fifth present more red flowers. From vertical view, the generated images also effectively capture and display the style of the reference images, especially the Ukiyo-e style.

### v2

@@ 5.2.2 Quantitative Comparison
Tab. 1 presents the style similarity, image quality, text alignment and the overall subjective preference of our method compared with the state-of-the-art methods. We can see that our method achieves the highest style similarity and content alignment, demonstrating that our method, through decoupling representations, effectively fixed the problem of semantic conflict and captured the overall artistic style. Furthermore, users demonstrate a significantly greater preference for our method over other ones. More quantitative results are provided in Supplementary Material Sec. 5.5.
5.3 Content Style Disentanglement
5.3.1 Qualitative illustration of disentanglement effect
@@
To verify the disentanglement effect, we conducted an experiment using portrait paintings as content and style respectively. As shown in Fig. 5, the generated images depict the content themes accurately: middle-aged men (first two rows), a young boy (third row), and middle-aged women (fourth and fifth rows). Vertically, the styles of the reference images are also well-preserved, with distinct brushstrokes in the third and fourth columns. This demonstrates the robustness of the proposed content-style disentanglement method. We also provided more disentanglement experiments using still life painting as content and style in Supplementary Material Sec. 7.
5.3.2 Impact of Content and Style Detail Levels on Disentanglement Performance
@@
From subfigure (a), we observe that as textual content descriptions increase in complexity, the generated images accurately reflect the described content. When an artistic image is used for content input, our method effectively disentangles content, capturing key elements like flowers, a vase, a table, and a teacup. Meanwhile, the generated images consistently replicate Monet’s brushstrokes and artistic style, regardless of of the level of detail in the content description.
@@
5.4 Ablation Study
To evaluate the contribution of each loss function in training the multimodal alignment and disentanglement network, we conducted an ablation study. The results are shown in Figure 7. ITC is the core loss function in the first stage. ITC ensures that the generated images maintain semantic similarity with the text input. Although the styles vary, all images consistently depict a village under a starry night, demonstrating ITC’s role in content control. ITG ensures that the model accurately extracts content information from the input image and utilizes it for text-guided generation, enabling more precise stylization. In Figure 7 (b), the generated images effectively incorporate Van Gogh’s brushstroke style in architectural structures, terrain, and the sky. ITM enhances the alignment between text and images (Figure 7 (c)), but at the cost of reduced stylization. Ultimately, these three loss functions work together to ensure that the model accurately represents the content described in the text while preserving the stylization effect, thereby improving text-guided artistic image generation (Figure 7 (d)). More ilustration for ablation study is provided in Supplementary Material Sec. 4.
6 Conclusion
In this paper, we proposed a multimodal approach for content and style representation disentanglement for artistic image stylization. We constructed a multimodal art dataset, WikiStyle+, to provide explicit supervision for decoupling. We employed contrastive learning tasks to learn disentangled content and style representations, which then guided a diffusion model to generate stylized images. Our experiments across various tasks demonstrated the superiority of our method and highlighted the importance of effective content and style disentanglement in image stylization.